Determine whether sqrt(49/100) is rational or irrational.
Determine whether -(16)^1/4 is rational or irrational. That is minus the 4th root of 16
Find the altitude of an isosceles triangle with sides of 13 inches and a base of 10 inches.
A right triangle has one side of 45 feet and the hypothenus is 51 feet. What is the length of the other side?

Answer by stanbon(75887)

Determine whether sqrt(49/100) is rational or irrational.
It's rational; it is 7/10.
------------------
Determine whether -(16)^1/4 is rational or irrational. That is minus the 4th root of 16
= -4
-----------------
Find the altitude of an isosceles triangle with sides of 13 inches and a base of 10 inches.
Draw the picture. The altitude forms two rt. triangles.
Use Pythagoras: altitude^2=13^2-5^2
altitude = sqrt 144 = 12 inches
--------------------
A right triangle has one side of 45 feet and the hypothenus is 51 feet. What is the length of the other side?
(other side)^2=51^2-45^2=576
other side = 24 feet.
Cheers,
Stan H.